### v1

Abstract
[…]
keywords:
linear layouts, recognition and characterization, priority queue layouts
category:
[…]
1 Introduction
Stack layouts were introduced by Bernhart and Kainen in 1979 [13] with the motivation to study a “quite natural” edge decomposition technique. A stack layout is a linear layout, that is, the vertices are positioned on a line $\ell$ with the edges being embedded on several half-planes, called pages, delimited by $\ell$. In a stack layout, edges embedded on the same page are forbidden to cross, a rule that contributed to their original designation as book embeddings. An example of stack layout on one page is depicted in Figure 1(b). The minimum number of pages required to obtain a stack layout of a graph is called the stack number. After their introduction, stack layouts attracted notable interest in the field of theoretical computer science due to their capacity to model certain aspects in several domains, including VLSI design [17, 18], fault-tolerant multiprocessing [18, 40], RNA folding [27], and traffic-light control [32]. This led, in particular, to a series of papers investigating upper bounds on the stack number of planar graphs [16, 28, 31] and culminating at an upper bound of four [42]. Bekos et al. [11] and Yannakakis [43] independently proved that this upper bound is tight.
The name stack layout derives from the fact that the edges assigned to the same page can be stored using a stack as follows [17]. Assume that $\ell$ is a horizontal line and consider the left-to-right ordering $v_{1},\ldots,v_{n}$ of the vertices. Now, consider a left-to-right sweep of $\ell$ and the set $E_{p}$ of edges assigned to a page $p$. Initially, the stack storing $E_{p}$ is empty. At vertex $v_{i}$, all edges that have $v_{i}$ as their right endpoint are popped from the stack. Since all vertices are at the same side of the half-plane delimited by $\ell$, the edges ending at $v_{i}$ are necessarily the last ones that have started prior to encountering $v_{i}$, so they are precisely the edges on top of the stack. Afterwards, edges starting at $v_{i}$ are pushed onto the stack. (Note that the order of the pushes is determined by the right endpoints of the corresponding edges.)
(a) Edge-weighted graph $G$.
(b) Stack layout of $G$.
(c) Queue layout of $G$.
(d) PQ-layout of $G$.
Figure 1: (a)–(c) Examples of stack and queue layouts on one page for the same graph $G$ ignoring edge weights. (d) A PQ-layout on one page of $G$ (edge weights are in blue).
Based on this latter interpretation of stack layouts, Heath, Leighton and Rosenberg [29, 30] proposed to study queue layouts of graphs where the edges occurring on the same page of the linear layout can be represented by a queue. More precisely, edges ending at $v_{i}$ are first dequeued from a queue, after which edges beginning at $v_{i}$ are enqueued during the left-to-right sweep of $\ell$. Thus, edges on the same page are allowed to cross but forbidden to properly nest. An example of a queue layout on one page is shown in Figure 1(c). Surprisingly, these quite non-planar representations still found many practical applications, including scheduling [14], VLSI [34], and, a decade after their inception, 3D crossing-free graph drawing [20, 23]. Similarly to stack layouts, one is interested in minimizing the number of pages required to obtain a queue layout, which is known as the queue number. However, in contrast to the stack number, the queue number of planar graphs remained elusive for several decades during which only sublinear (but non-constant) upper bounds have been found [1, 5, 6, 21]. The first significant difference between planar and non-planar graphs was only demonstrated in 2019 when it was shown that bounded degree planar graphs have bounded queue number [9], whereas non-planar bounded degree graphs were already known to not have this property [41]. Recently, Dujmović et al. [22] proved that the queue number of planar graphs is bounded by a constant. Minor improvements have since been made [10, 25].
Another interesting recent direction in investigating linear layouts is to expand beyond stack and queue layouts by using other data structures for partitioning edges. For instance, mixed linear layouts allow the usage of both stack and queue pages [2, 19, 37], whereas deque layouts use double-ended queues which allow to insert and pull edges from the head and the tail [3, 4, 12], and rique layouts use restricted-input queues which are double-ended queues where insertions are allowed to occur only at the head [8].
Our contribution.
The wide range of applications of linear layouts and the fact that they have been restricted so far to unweighted graphs, naturally motivate extensions of the notion of linear layout to edge-weighted graphs. For example, in scheduling applications, the edges of a graph can represent processes, each with an associated priority; each vertex $v$ of the graph enforces that all processes corresponding to edges incident to $v$ begin simultaneously; among the running processes, those with higher priority must be completed first. Edge weights can also model constraints in other classical applications of linear layouts, such as logistics networks modelled as switchyard networks [33, 36], where adjacent edges represent containers that must be pushed or popped simultaneously across different storage locations. The weight of an edge corresponds to the weight of its container, and it may be necessary to ensure that containers stacked on top of others are lighter.
We introduce a model called priority queue layout, or simply PQ-layout. This model utilizes priority queues for the storage of edges assigned to the same page, where the edge weights correspond to the priorities (i.e., keys) for the data structure. As for stack and queue layouts, in a PQ-layout all vertices of the graph lie on a horizontal line $\ell$, and the edges are assigned to different priority queues according to a left-to-right sweep of $\ell$. When an edge $e$ is dequeued, it must have the minimum priority among the edges stored in the same priority queue as $e$. See Figure 1(d) for an example, and refer to Section 2 for a formal definition of our model. Note that, unlike stack and queue layouts, our model allows edges in the same page to cross or properly nest, provided that they have suitable weights.
Following a current focus of research on linear layouts, we mainly study the priority queue number of edge-weighted graphs, that is, the number of pages required by their PQ-layouts:
[…]
We remark that, as far as we know, there is only one previous study in the literature about using linear layouts of edge-weighted graphs [7]; however, this study uses edge weights to further restrict stack layouts. For space restrictions, proofs of statements marked with a (clickable) ($\star$) are omitted or sketched. See the appendix for a full version of these proofs.
2 PQ-Layouts
Let $\langle G,w\rangle$ be an edge-weighted graph, that is, $G=(V,E)$ is an undirected graph and $w\colon E\to\mathbb{R}$ is an edge-weight function. A priority queue layout, or simply a PQ-layout, $\Gamma$ of $\langle G,w\rangle$ consists of a linear (left-to-right) ordering $v_{1}\prec\cdots\prec v_{n}$ of the vertices of $G$ and a partitioning of the edges of $G$ into $k$ sets $\mathcal{P}_{1},\ldots,\mathcal{P}_{k}$, where each set is called a page and is associated with a priority queue. For each page $\mathcal{P}_{i}$ with $i\in\{1,\dots,k\}$, the following must hold: when traversing the linear ordering $\prec$ (from left to right) and performing, at each encountered vertex $v\in V$, the operations O.1 and O.2 (in this order), condition C.1 must always hold.
  1. O.1
    Each edge $e\in\mathcal{P}_{i}$ having $v$ as its right endpoint is pulled from the priority queue. We do not count $e$ as active any more.
  2. O.2
    Each edge $e\in\mathcal{P}_{i}$ having $v$ as its left endpoint is inserted into the priority queue of $\mathcal{P}_{i}$ with key $w(e)$. We say that $e$ becomes active. Two edges being active at the same time (even though they may have become active in different steps) are called co-active.
  3. C.1
    There is no pair of co-active edges $e$ and $e^{\prime}$ such that $e$ has $v$ as its right endpoint, $e^{\prime}$ has a distinct right endpoint, and $w(e)>w(e^{\prime})$. In other words, among the active edges of $\mathcal{P}_{i}$, the edges having $v$ as their right endpoint have the smallest edge weights.
PQ-layouts can be equivalently defined by the absence of the forbidden configuration F.1.
  1. F.1
    For some $i\in\{1,\dots,k\}$, there are two edges $e=uv$ and $e^{\prime}=u^{\prime}v^{\prime}$ such that $e,e^{\prime}\in\mathcal{P}_{i}$, $w(e)>w(e^{\prime})$, and $u,u^{\prime}\prec v\prec v^{\prime}$.
The order of $u$ and $u^{\prime}$ is irrelevant for F.1. For completeness, the resulting three arrangements of the involved vertices are illustrated in Figures 2(a), 2(b) and 2(c). Note that the forbidden pseudo-nesting in Figure 2(b) requires $u=u^{\prime}$. In contrast, the symmetric case where $v=v^{\prime}$ is never forbidden. This is a direct consequence of the asymmetry between the insert and pull operations of priority queues. Thus, unlike for stack and queue layouts, the reverse vertex ordering does not necessarily provide a PQ-layout.
(a) Forbidden
(a) Nesting
(b) Forbidden
(b) Pseudo-Nesting
(c) Forbidden
(c) Crossing
(d) $k$-Inversion
Figure 2: (a)–(c) Arrangements in a PQ-layout that correspond to the Forbidden Configuration F.1. (d) A $k$-inversion, a forbidden arrangement in linear layouts with priority queue number $k-1$.
We refer to the linear ordering $\prec$ of vertices also by the ordering along the spine. For a PQ-layout $\Gamma$ of any graph, let $\operatorname{pqn}(\Gamma)$ denote the number of priority queues in $\Gamma$. Now, the priority queue number $\operatorname{pqn}(G,w)$ of $\langle G,w\rangle$ is the minimum integer $k\geq 0$ for which there exists a PQ-layout $\Gamma$ of $\langle G,w\rangle$ with $\operatorname{pqn}(\Gamma)=k$.
For any given graph $G$, there always exists an edge-weight function $w\colon E\to\mathbb{R}$ such that $\operatorname{pqn}(G,w)=1$. It is enough to assign the same weights to all edges but it is also possible to use distinct edge weights. Namely, define any linear ordering of the vertices $v_{1}\prec\cdots\prec v_{n}$ and visit the vertices in this ordering. When a vertex $v$ with $v\neq v_{1}$ is visited, assign consecutive (non-negative) integer weights to all edges for which $v$ is the right endpoint such that the weights monotonically increase. Hence, any graph with any given linear ordering of its vertices can have priority queue number 1 if we can choose the edge-weight function. However, the edge-weight function may be provided as part of the input. Thus, it is interesting to study bounds on the priority queue number of families of graphs that hold independently of the edge-weight function. To this end, for a graph $G=(V,E)$ without any prescribed edge-weight function, the priority queue number $\operatorname{pqn}(G)$ is the minimum integer $k\geq 0$ such that for every possible edge-weight function $w\colon E\to\mathbb{R}$ we have $\operatorname{pqn}(G,w)\leq k$.
Some of our proofs utilize configurations where every pair of edges must be assigned to distinct priority queues. Let $\langle G,w\rangle$ be an edge-weighted graph, let $\prec$ be a vertex ordering of $G$, and let $k$ be an integer such that $k\geq 1$. We define a $k$-inversion to be a sequence $(e_{1},\ldots,e_{k})$ of $k$ edges in $G$ with pairwise distinct right endpoints $v_{1},\ldots,v_{k}$, respectively, such that (i) $v_{k}\prec\cdots\prec v_{1}$, where $v_{k}$ and $v_{1}$ are called the left and right end, respectively, (ii) the left endpoint of each of $e_{1},\ldots,e_{k}$ appears to the left of $v_{k}$ in $\prec$, and (iii) $w(e_{1})<\cdots<w(e_{k})$. See Figure 2(d) for an illustration. Any two edges in a $k$-inversion form a forbidden configuration and thus must be assigned to distinct priority queues. In other words, if, for edge weights $w$ of $G$, every vertex ordering of $\langle G,w\rangle$ contains a $k$-inversion, then $\operatorname{pqn}(G,w)\geq k$ and hence $\operatorname{pqn}(G)\geq k$.
3 Priority Queue Number of Complete Graphs
In this section, we concentrate on complete graphs and complete bipartite graphs. We start with a simple upper bound for the complete graph $K_{n}$ on $n$ vertices.
{observation}
$\operatorname{pqn}(K_{n})\leq n-1$.
Proof.
Choose an arbitrary ordering of the vertices $v_{1}\prec\cdots\prec v_{n}$ and define the sets (priority queues) $\mathcal{P}_{1},\dots,\mathcal{P}_{n-1}$. For each $j\in\{2,\ldots,n\}$, assign all edges whose right endpoint is $v_{j}$ to $\mathcal{P}_{j-1}$. Clearly, no forbidden configuration occurs, independent of the edge-weight function. ∎
Next, we consider bipartite graphs, that is, the graphs whose vertex sets are the union of two disjoint independent sets $A$ and $B$. In the literature on linear layouts, separated linear layouts, where vertices of $A$ precede the vertices of $B$ or vice versa, have been considered for bipartite graphs [24]. The separated priority queue number $\operatorname{spqn}(G)$ of a bipartite graph $G$ is the minimum $k$ such that, for any edge-weight function, there is a PQ-layout $\Gamma$ that is a separated linear layout and $\operatorname{pqn}(\Gamma)=k$. We can easily determine the separated priority queue number of the complete bipartite graph $K_{n,n}$. At the end of this section, we will use the separated priority queue number to give a lower bound on the priority queue number of complete and complete bipartite graphs.
Theorem 3.1.
$\operatorname{spqn}(K_{m,n})=\min\{m,n\}$.
Proof 3.2.
First observe that $\operatorname{spqn}(K_{m,n})\leq\min\{m,n\}$: place the vertices of the larger set, say $A$, before the vertices of the smaller set, say $B$. Then, independent of the edge weight function, for each vertex $b\in B$, all edges ending at $b$ can be assigned to a separate page.
To show $\operatorname{spqn}(K_{m,n})\geq\min\{m,n\}$, assume w.l.o.g. that the vertices in $A$ precede the vertices in $B$. Consider an edge weight function where, at every vertex $b\in B$, each edge weight in $\{1,2,\dots,\min\{m,n\}\}$ occurs at least once among $b$’s incident edges. In any bipartite PQ-layout $\Gamma$ of $K_{m,n}$, denote the vertices of $B$ by $b_{1},b_{2},\dots$ in the order they appear along the spine. In $\Gamma$, $b_{1}$ is incident to an edge with weight $\min\{m,n\}$, $b_{2}$ is incident to an edge with weight $\min\{m,n\}-1$, etc., $b_{\min\{m,n\}}$ is incident to an edge of weight 1. This is a $\min\{m,n\}$-inversion requiring at least $\min\{m,n\}$ pages.
Theorem 3.1 also provides an upper bound for the (non-separated) priority queue number of $K_{n,n}$. Next, we give a corresponding linear lower bound showing that $\operatorname{pqn}(K_{n,n})\in\Theta(n)$.
[…]
$\operatorname{pqn}(K_{n,n})\geq\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil\approx 0.191n$.
Proof 3.4 (Proof Sketch).
For $K_{n,n}$, we define an edge-weight function $w$ that has, at each vertex, for every $i\in\{1,\ldots,n\}$, exactly one incident edge of weight $i$ [^1] In contrast to the proof of Theorem 3.1, we cannot arbitrarily assign weights to edges incident to each vertex $b\in B$ because this may give twice the same edge weight at a vertex of partition $A$. To this end, we partition the edge set into $n$ perfect matchings $M_{1},\ldots,M_{n}$, and, for each edge $e\in M_{i}$, we set $w(e)=i$.
For any PQ-layout $\Gamma$ of $\langle K_{n,n},w\rangle$, we can find a sublayout $\Gamma^{\prime}$ that is a bipartite PQ-layout of $\langle K_{\frac{n}{2},\frac{n}{2}},w\rangle$ – either take the first $\frac{n}{2}$ vertices of $A$ and the last $\frac{n}{2}$ vertices of $B$ or vice versa. [^2] For simplicity, we assume that $n$ is even.
Note that we cannot directly apply Theorem 3.1 to $\Gamma^{\prime}$ because we have already fixed an edge-weight function. Instead, we analyze the possible distribution of the edge weights incident to vertices in $\Gamma^{\prime}$ by a representation as a black-and-white grid; see the appendix for an illustration and a more extensive description. [^3] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. Our grid has $\frac{n}{2}$ columns that represent the vertices of the latter partition (in order from left to right as they appear in $\Gamma^{\prime}$) and it has $n$ rows that represent the edge weights $\{1,\dots,n\}$ (in order from bottom to top). A grid cell in column $i$ and row $j$ is colored black if the $i$-th vertex is incident to an edge of weight $j$, and it is colored white otherwise. Observe that, in this grid, a strictly monotonically decreasing path of black cells having length $k$ is equivalent to a $k$-inversion in $\Gamma^{\prime}$. We can show that there always exists such a path of length $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$: the candidates for being the first black cell of the path are those who do not have another black cell in their top left region of the grid. We find these cells, color them white and iteratively repeat this process to find the candidates for being the second, third, etc. black cell of the path. We can prove that we need at least $\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$ iterations until the grid is white.
Since $K_{n}$ contains $K_{\lfloor\frac{n}{2}\rfloor,\lfloor\frac{n}{2}\rfloor}$ as a subgraph, Theorem 3.3 immediately implies a linear lower bound on $\operatorname{pqn}(K_{n})$:
Corollary 3.5.
$\operatorname{pqn}(K_{n})\geq\left\lfloor\frac{3-\sqrt{5}}{8}n\right\rfloor\approx 0.0955n$.
4 Characterizing Graphs with Priority Queue Number 1
We study which graphs have priority queue number 1. We will provide a characterization of the graphs with priority queue number 1 in terms of forbidden minors (Section 4.4). We first show that the family of graphs with priority queue number 1 is minor-closed (Section 4.1). Then, we describe families of graphs with priority queue number 1 (Section 4.2) and a set of forbidden minors (Section 4.3). The characterization of Section 4.4 follows by proving that every graph that does not contain any of the forbidden minors of Section 4.3 falls in one of the families of Section 4.2. This characterization leads to an efficient recognition algorithm.
4.1 PQ-Layouts of Minors
A minor $G^{\prime}$ of a graph $G=(V,E)$ is obtained from $G$ by a series of edge contractions and by removing vertices and edges. Formally, for an edge $e=uv$ in $G$, the contraction of edge $e$ yields the graph $G/e$, obtained by replacing vertices $u$ and $v$ by a single vertex $x_{uv}$ with incident edges $\{x_{uv}y\mid uy\in E\text{ or }vy\in E\}$. Many important graph classes (e.g., planar graphs) are minor-closed, that is, if we take a minor of a graph lying in such a class, the resulting graph belongs to this class as well. Every minor-closed graph class is defined by a finite set of forbidden minors and can be recognized efficiently [38, 39].
[…]
Let $G=(V,E)$ be a graph with $\operatorname{pqn}(G)=1$ and let $G^{\prime}=G/e_{0}$ be obtained from $G$ by contracting an edge $e_{0}=uv\in E$. Then, $\operatorname{pqn}(G^{\prime})=1$.
Proof 4.2 (Proof Sketch).
We prove $\operatorname{pqn}(G^{\prime},w^{\prime})=1$ for any given edge-weight function $w^{\prime}$. We construct a PQ-layout $\Gamma$ of $\langle G,w\rangle$ using one priority queue where $w(e)=w^{\prime}(e)$ except for $w(e_{0})=W<\min_{e\in E\setminus\{e_{0}\}}w(e)$. Assuming $u\prec v$ in $\Gamma$, we construct a PQ-layout $\Gamma^{\prime}$ of $\langle G^{\prime},w^{\prime}\rangle$: We place vertex $x_{uv}$ obtained from contracting $e_{0}$ at the position of $u$ in $\Gamma$. Vertices preceding $u$ or succeeding $v$ in $\Gamma$ retain their positions in $\Gamma^{\prime}$, and vertices between $u$ and $v$ in $\Gamma$ are sorted according to the weight of their edges shared with $x_{uv}$. See Appendix B for details.
Corollary 4.3.
The family of graphs with priority queue number 1 is minor-closed.
4.2 Graphs with Priority Queue Number 1
We begin by showing how to construct PQ-layouts of trees:
Lemma 4.4.
Let $\langle T=(V,E),w\rangle$ be an $n$-vertex edge-weighted tree with root $r\in V$. Then, a PQ-layout $\Gamma$ of $\langle T,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$, where for each vertex $v\neq r$ its parent $p(v)$ precedes $v$ in the linear ordering $\prec$ along the spine, can be constructed in $\mathcal{O}(n\log n)$ time.
Proof 4.5.
We describe an algorithm that constructs a PQ-layout $\Gamma$ of $T$; see Figure 3 for an example. For each vertex $v\in V\setminus\{r\}$, we define its weight $w(v)$ as $w(v)=w(vp(v))$. As an initial layout, we first place $r$ as the leftmost vertex and the children of $r$ in increasing order of their weights to the right of $r$. During our algorithm, we keep track of the last expanded vertex $v^{\star}$, which is the rightmost vertex whose children (if any) have already been placed. Note that initially $v^{\star}=r$. We now maintain the following invariants:
  1. T.1
    For each vertex $v$ with $v\preceq v^{\star}$, all children have already been placed.
  2. T.2
    For each vertex $v$ with $v^{\star}\prec v$, no child has been placed yet. In addition, $p(v)\preceq v^{\star}$.
  3. T.3
    Let $S$ denote the set of already placed vertices succeeding $v^{\star}$ in $\prec$. Then, the vertices in $S$ occur in $\prec$ in increasing order of their weights.
  4. T.4
    The already constructed PQ-layout has priority queue number 1.
It is easy to see that the invariants hold for the initial layout. We now iteratively consider the vertex $v^{\prime}$ immediately succeeding $v^{\star}$ in $\prec$ and perform the following operations. For each child $c$ of $v^{\prime}$, we place $c$ to the right of $v^{\prime}$ such that $S\cup\{c\}\setminus\{v^{\prime}\}$ is ordered by weight. This can be easily done since by T.3, $S$ (and thus also $S\setminus\{v^{\prime}\}$) is already ordered by weight, i.e., there is a unique position for each $c$ and T.3 holds again after $v^{\prime}$ has become the new $v^{\star}$. Next, consider edge $v^{\prime}c$. By T.4, the only way that the resulting linear layout has not priority queue number 1 is if $v^{\prime}c$ is included in a forbidden configuration. Consider an edge $uv$ with $u\prec v$ that is co-active with $v^{\prime}c$. Clearly, $v\in S$ and by T.2, we have that $u\preceq v^{\prime}$ and $u=p(v)$. By T.3, it follows that $c$ and $v$ are sorted by their weights, which are equal to $w(v^{\prime}c)$ and $w(uv)$, thus, $uv$ and $v^{\prime}c$ form no forbidden configuration, i.e., T.4 is maintained.
(a) A weighted rooted tree $T$
(b) A PQ-layout $\Gamma$ of $T$.
Figure 3: A PQ-layout of a weighted rooted tree computed with the algorithm in Lemma 4.4.
After placing all children of $v^{\prime}$, we set $v^{\star}=v^{\prime}$. Clearly, T.1 is maintained as we placed all children of $v^{\prime}$ whereas by induction the children of the vertices preceding $v^{\prime}$ have already been placed. Moreover, T.2 is maintained as we only placed all children of $v^{\prime}$ to the right of $v^{\prime}$ whereas by induction the children of other vertices of $S$ are not placed yet.
For every vertex, we need to find its position depending on the weight. This can be done in $\mathcal{O}(\log n)$ time if we maintain a suitable search data structure. The rest of the insertion can be done in constant time. Overall, this results in a running time in $\mathcal{O}(n\log n)$.
Next, we construct a specific layout of caterpillars, which are a special subclass of trees. Namely, a caterpillar $C$ consists of an underlying path $P(C)$ and additional leaves, each having exactly one neighbor on $P(C)$. Be aware that, for the same caterpillar $C$, there are several choices for $P(C)$, and in particular $P(C)$ might end in a degree-$1$ vertex of $C$.
Lemma 4.6.
Let $\langle C,w\rangle$ be an $n$-vertex edge-weighted caterpillar with underlying path $P(C)$ and let $r$ be one of the two degree-$1$ vertices in $P(C)$. Then, a PQ-layout $\Gamma$ of $\langle C,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ where $r$ is the rightmost vertex can be constructed in $\mathcal{O}(n)$ time.
Proof 4.7.
We label $P(C)$ as $p_{1}p_{2}{\ldots}p_{k}=r$ and arrange the vertices of $P(C)$ in the order $p_{1}\prec p_{2}\prec\ldots\prec r$. Then for $i\in\{1,\ldots,k\}$, we place all leaves at $p_{i}$ between $p_{i-1}$ and $p_{i}$ in an arbitrary order (note that for $p_{1}$, we place its leaves before $p_{1}$). It is easy to see that the resulting layout contains no forbidden configuration and the statement follows.
We now shift our attention to graphs containing cycles. First, we consider cycles alone.
Lemma 4.8.
Let $\langle C=(V,E),w\rangle$ be an $n$-vertex edge-weighted cycle, and let $v$ be any given vertex of $V$. Then, a PQ-layout $\Gamma$ of $\langle C,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ where $v$ is the leftmost vertex can be constructed in $\mathcal{O}(n)$ time.
Proof 4.9.
We maintain two candidate vertices $a$ and $b$ to be placed next. We call the neighbor of $a$ ($b$, resp.) that has already been placed anchor vertex $a^{\prime}$ ($b^{\prime}$, resp.). Initially, $a$ and $b$ are the two neighbors of the leftmost vertex $v$ and $v=a^{\prime}=b^{\prime}$. We iteratively append the candidate vertex whose edge to its anchor vertex has the smallest weight to the right of the linear order (if both weights are equal, we take any of both candidates). If $w(a^{\prime}a)\leq w(b^{\prime}b)$, we append $a$ while we do not yet place $b$. After appending $a$, we set $a^{\prime}=a$ and the other neighbor of $a$ becomes the new candidate vertex $a$. We proceed symmetrically with $b$ if $w(a^{\prime}a)>w(b^{\prime}b)$. When $a=b$, we place the last remaining vertex $a$ as the rightmost vertex.
Clearly, every vertex is placed. We show that no forbidden configuration occurs. Suppose that there are two co-active edges $uv$ (where $u\prec v$) and $xy$ (where $x\prec y$) with $w(uv)>w(xy)$ and $v\prec y$. In the course of our algorithm, after $x$ and $u$ have been placed and before $v$ and $y$ have been placed, the candidate vertices were $v$ and $y$ whose anchor vertices were $u$ and $x$. There could not have been a different anchor or candidate vertex because $C$ is a cycle. Then, however, we would not have placed $v$ first because $w(uv)>w(xy)$ – a contradiction.
Note that the running time of the initialization, termination and each iterative step is constant. Therefore, the overall running time is in $\mathcal{O}(n)$.
We then consider the family of legged cycles. Namely, a legged cycle $L$ consists of a single underlying cycle $C(L)$ and additional leaves, each having exactly one neighbor on $C(L)$.
[…]
Let $\langle L,w\rangle$ be an $n$-vertex edge-weighted legged cycle with underlying cycle $C(L)$. Then, a PQ-layout $\Gamma$ of $\langle L,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ can be constructed in $\mathcal{O}(n\log n)$ time.
Proof 4.11 (Proof Sketch).
We remove the heaviest edge $e^{\star}$ of $C(L)$ and all leaves where the weight of the incident edge is greater than $w(e^{\star})$. We lay out the remaining graph using Lemma 4.6 such that the endpoints of $e^{\star}$ are the first and the last vertex. We reinsert $e^{\star}$ and, in increasing order of the weights of their incident edges, we add the initially removed leaves to the right.
Next, we show that we may attach one caterpillar (Lemma 4.12) or two caterpillars to a single quadrangle (Lemma 4.14) or a single triangle (Corollary 4.16).
Lemma 4.12.
Let $\langle G,w\rangle$ be an $n$-vertex edge-weighted graph consisting of a cycle $O$ and a caterpillar $C$ with underlying path $P(C)$ such that $V(O)\cap V(C)=\{r\}$ is a degree-$1$ vertex of $P(C)$. Then, a PQ-layout $\Gamma$ of $\langle G,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ can be constructed in $\mathcal{O}(n)$ time.
Proof 4.13.
We first draw $C$ and $O$ separately using Lemmas 4.6 and 4.8, respectively. Then, we identify the two occurrences of $r$ in $C$ and $O$ with each other, which yields our PQ-layout $\Gamma$ of $G$. This does not result in a forbidden configuration since $r$ is the rightmost vertex of $C$ and the leftmost vertex of $O$ and, hence, there is no edge spanning over $r$.
[…]
Let $\langle G,w\rangle$ be an $n$-vertex edge-weighted graph consisting of:
  • A $4$-cycle $\Box=abcd$.
  • A caterpillar $C_{a}$ with underyling path $P(C_{a})$ starting at vertex $a\in\Box$.
  • A caterpillar $C_{c}$ with underyling path $P(C_{c})$ starting at vertex $c\in\Box$.
Then, a PQ-layout $\Gamma$ of $\langle G,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ can be constructed in $\mathcal{O}(n\log n)$.
Proof 4.15 (Proof Sketch).
We arrange the cycle such that only the heaviest edge is co-active with the other three edges. We use Lemma 4.4 for laying out the one caterpillar with designated leftmost vertex and we use Lemma 4.6 for laying out the other caterpillar with designated rightmost vertex. In the appendix, we argue that we can combine these three subgraphs, potentially adding some leaves later on, while still using only one priority queue.
By Lemma 4.1, if $G$ is a minor of a graph as described in Lemma 4.14, then also $\operatorname{pqn}(G)=1$.
Corollary 4.16.
If $G$ consists of a triangle $\triangle=abc$ and two caterpillars $C_{a}$ and $C_{b}$ with underlying paths starting at vertex $a\in\triangle$ and $b\in\triangle$, respectively, then $\operatorname{pqn}(G)=1$.
Finally, we consider graphs that contain more than one cycle. If we exclude disconnected graphs, there exist exactly two such graphs with priority queue number 1, the complete bipartite graph $K_{2,3}$ and the graph obtained from $K_{4}$ by removing an edge. We will later see that these are the only ones.
[…]
$\operatorname{pqn}(K_{2,3})=1$.
In addition, by Lemma 4.1, we obtain the following from Lemma 4.17.
Corollary 4.18.
$\operatorname{pqn}(K_{4}-e)=1$ where $K_{4}-e$ is obtained from $K_{4}$ by removing one edge.
4.3 Forbidden Minors for Graphs with Priority Queue Number 1
Each of the graphs $\mathcal{F}_{1},\ldots,\mathcal{F}_{8}$ shown in Figure 4 has priority queue number at least two. The numbers at the edges specify edge-weight functions for which one priority queue does not suffice. In the appendix, we prove the following observation:
[…]
In any PQ-layout $\Gamma$ of an edge-weighted cycle $\langle C,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$, the last vertex on the spine is incident to an edge of maximum weight in $\langle C,w\rangle$.
We now use Lemma 4.19 to prove the main result of this subsection:
[…]
For each graph $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$ (see Figure 4), $\operatorname{pqn}(\mathcal{F})>1$ holds.
Proof 4.21 (Proof of Cases $\mathcal{F}_{1}$$\mathcal{F}_{3}$).
To show $\operatorname{pqn}(\mathcal{F})>1$, it suffices to show $\operatorname{pqn}(\mathcal{F},w)>1$ for some edge weighting $w$. We claim that the edge weightings in Figure 4 require more than one priority queue. We consider each of the eight graphs individually and suppose for a contradiction that there is a PQ-layout $\Gamma$ with $\operatorname{pqn}(\Gamma)=1$. We consider $\mathcal{F}_{1}$$\mathcal{F}_{3}$ here and $\mathcal{F}_{4}$$\mathcal{F}_{8}$ in the appendix.
(a) $\mathcal{F}_{1}$
(b) $\mathcal{F}_{2}$
(c) $\mathcal{F}_{3}$
(d) $\mathcal{F}_{4}$
(e) $\mathcal{F}_{5}$
(f) $\mathcal{F}_{6}$
(g) $\mathcal{F}_{7}$
(h) $\mathcal{F}_{8}$
Figure 4: Forbidden minors for graphs with priority queue number 1. The graphs with the provided edge weights do not admit a linear layout with priority queue number 1. The thickness of an edge indicates the weight of that edge.
$\mathcal{F}_{1}$
: Consider the triangle $\triangle abd$. By Lemma 4.19, $a$ or $b$ is the last vertex of $\triangle abd$ (on the spine) since $ab$ is its heaviest edge. In the triangle $\triangle bcd$, $cd$ is the heaviest edge, which implies that $c$ or $d$ is the last vertex of $\triangle bcd$. Hence, neither $b$ nor $d$ can be the (overall) rightmost vertex. Since the heaviest edge of the 4-cycle $abcd$ is again $ab$, $a$ is the rightmost vertex. In the triangle $\triangle acd$, $cd$ is the heaviest edge, which implies that $a$ cannot be the rightmost vertex; a contradiction.
$\mathcal{F}_{2}$
: By Lemma 4.19 and symmetry, we may assume that $a$ and $d$ are the last vertices of $\triangle abc$ and $\triangle cde$, respectively. Moreover, assume by symmetry that $a\prec d$, i.e., $c\prec a\prec d$. However, then edge $ab$ ends below the lighter edge $cd$; a contradiction.
$\mathcal{F}_{3}$
: By Lemma 4.19, $c$ or $d$ is the last vertex of the 5-cycle $abcde$. In the 4-cycle $bcde$, however, $b$ or $e$ is the last vertex; a contradiction.
4.4 Characterization and Recognition
We will now combine the results shown in the previous subsections to show the following:
[…]
Let $G$ be a graph. Then, $\operatorname{pqn}(G)=1$ if and only if $G$ does not contain any $\mathcal{F}_{i}$ for $i\in\{1,\ldots,8\}$ as a minor.
Proof 4.23 (Proof Sketch).
We analyze the structure of $G$ and we see that either $G$ is a graph that has $\operatorname{pqn}(G)=1$ due to a result from Section 4.2, or $G$ has some $\mathcal{F}_{i}$ for $i\in\{1,\ldots,8\}$ as a minor.
If $G$ is a tree, then $\operatorname{pqn}(G)=1$ by Lemma 4.4. If $G$ contains exactly one cycle $C$, then $G-E(C)$ is a forest $F$, and we consider each component of $F$ as rooted at $C$. If every component is an isolated vertex or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma 4.10. If some component is more complex than a rooted star or a rooted caterpillar, then $G$ contains $\mathcal{F}_{4}$ as a minor. Now suppose some component $K$ is a rooted caterpillar. If all other components are isolated vertices, then $\operatorname{pqn}(G)=1$ by Lemma 4.12. If there are two further non-trivial components, then $G$ contains $\mathcal{F}_{5}$ as a minor. So assume that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{8}$ as a minor. If none of the above applies, either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Corollary 4.16, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma 4.14. The case where $G$ contains at least two cycles remains. If $G$ has two edge-disjoint cycles, then $G$ contains $\mathcal{F}_{2}$ as a minor. If two cycles have at least two vertices but no edge in common, then $G$ contains $\mathcal{F}_{7}$ as a minor. If any two cycles share at least one edge, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas 4.17 and 4.18, or $G$ contains $\mathcal{F}_{1}$, $\mathcal{F}_{6}$, or $\mathcal{F}_{7}$ as a minor.
Theorem 4.22 implies that the structure of a graph $G$ with $\operatorname{pqn}(G)=1$ is quite limited. In fact, forbidden minors $\mathcal{F}_{1},\mathcal{F}_{2}$ and $\mathcal{F}_{3}$ imply that $G$ can contain at most three cycles. More precisely, if $G$ contains two cycles, it is necessarily a $K_{2,3}$ or a $K_{4}-e$ since forbidden minors $\mathcal{F}_{6}$ and $\mathcal{F}_{7}$ forbid any other edge. Moreover, if $G$ contains a single cycle, forbidden minor $\mathcal{F}_{4}$ implies that no non-trivial tree (i.e., a tree that is more than a caterpillar) can be attached to the cycle. Finally, forbidden minors $\mathcal{F}_{5}$ and $\mathcal{F}_{8}$ dictate how legs can be combined with an attached caterpillar. $\mathcal{F}_{5}$ implies that at most one vertex of the cycle can have legs whereas $\mathcal{F}_{8}$ implies that, for a cycle of length $4$, legs can only be attached to the vertex opposite of the vertex attached to the caterpillar whereas for cycles of lengths larger than $4$, no legs are allowed. These observations in conjunction with Lemmas 4.4, 4.6, 4.10, 4.14, 4.16, 4.17 and 4.18 imply the following:
Corollary 4.24.
Given a graph $G$, it can be decided in $\mathcal{O}(n)$ time if $\operatorname{pqn}(G)=1$. Moreover, if $\operatorname{pqn}(G)=1$, a PQ-layout $\Gamma$ of $\langle G,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ can be computed in $\mathcal{O}(n\log n)$ time for a given edge-weight function $w$.
5 PQ-Layouts of Graphs with Bounded Pathwidth and Treewidth
We present results for graphs of bounded pathwidth and bounded treewidth. We assume familiarity with these concepts; otherwise see Appendix C for formal definitions.
Theorem 5.1.
Let $G$ be a graph with pathwidth at most $p$. Then, $\operatorname{pqn}(G)\leq p+1$.
Proof 5.2.
We may assume w.l.o.g. that $G=(V,E)$ is edge-maximal of pathwidth $p$, i.e., $G$ is an interval graph with clique number $\omega(G)=p+1$ [15]. Let $\{I_{v}=[a_{v},b_{v}]\}_{v\in V}$ be an interval representation of $G$ with distinct interval endpoints. In particular $b_{u}\neq b_{v}$ for any $u\neq v\in V$.
We take the ordering $v_{1},\ldots,v_{n}$ of vertices given by their increasing right interval endpoints. That is, $b_{v_{1}}<\cdots<b_{v_{n}}$. Crucially, for any $i<j<k$ with $v_{i}v_{k}\in E$, also $v_{j}v_{k}\in E$.
To define the partition of $E$ into $p+1$ priority queues $\mathcal{P}_{1},\ldots,\mathcal{P}_{p+1}$, we consider a proper vertex coloring of $G$ with $\chi(G)=\omega(G)=p+1$ colors. This exists, as $G$ is an interval graph. Now for $c=1,\ldots,p+1$ let $\mathcal{P}_{c}$ be the set of all edges in $G$ whose right endpoint in the vertex ordering has color $c$. In fact, each $\mathcal{P}_{c}$ contains none of the forbidden configurations in Figure 2, and hence is a priority queue independent of the edge-weights, since in each forbidden configuration in Figure 2 there would be an edge in $G$ between the right endpoints $v_{r}$ and $v_{r^{\prime}}$ of the two edges $e$ and $e^{\prime}$ forming the forbidden configuration. But this would imply different colors for $v_{r}$ and $v_{r^{\prime}}$ and therefore different priority queues for $e$ and $e^{\prime}$.
We remark that Corollary 3.5 implies that there are graphs with pathwidth $p$ and priority queue number at least $\left\lfloor\frac{3-\sqrt{5}}{8}(p+1)\right\rfloor$ as $K_{p+1}$ has pathwidth $p$.
Next, we shall construct for every integer $p\geq 1$ a graph $G=G_{p}$ of treewidth $2$ together with edge weights $w$ so that every vertex ordering $\sigma$ of $G$ contains a $p$-inversion. If every vertex ordering of $\langle G,w\rangle$ contains a $p$-inversion, then $\operatorname{pqn}(G,w)\geq p$ and hence $\operatorname{pqn}(G)\geq p$.
[…]
For every $p$ there is a graph $G_{p}$ of treewidth $2$ and $\operatorname{pqn}(G_{p})\geq p$.
Proof 5.4 (Proof Sketch).
Let $p\geq 1$ be a fixed integer. Our desired graph $G_{p}$ will be a rooted $2$-tree, i.e., an edge-maximal graph of treewidth $2$ with designated root edge, which are defined inductively through the following construction sequence:
  • A single edge $e^{*}=v^{\prime}v^{\prime\prime}$ is a $2$-tree rooted at $e^{*}$.
  • If $G$ is a $2$-tree rooted at $e^{*}$ and $e=uv$ is any edge of $G$, then adding a new vertex $t$ with neighbors $u$ and $v$ is again a $2$-tree rooted at $e^{*}$. We say that $t$ is stacked onto $uv$.
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $v^{\prime}$, $v^{\prime\prime}$) appears to the right of both its parents. First we show that it is enough to force a $p$-inversion (and hence $\operatorname{pqn}(G)\geq p$) in every left-growing vertex ordering.
[…]
If $\langle G,w\rangle$ is an edge-weighted rooted $2$-tree so that every left-growing vertex ordering contains a $p$-inversion, then there exists an edge-weighted rooted $2$-tree $\langle\tilde{G},\tilde{w}\rangle$ so that every vertex ordering contains a $p$-inversion.
{claimproof}
[Proof Sketch] Intuitively, whenever in the construction sequence of $G$ a vertex $t$ is stacked onto an edge $uv$, in $\tilde{G}$ we instead stack $p^{2}$ “copies” of $t$ called $t_{1},\ldots,t_{p^{2}}$ onto $uv$ (treating each $t_{i}$ like $t$ henceforth). The weights $\tilde{w}(ut_{i})$ and $\tilde{w}(vt_{i})$, $i=1,\ldots,p^{2}$, are chosen very close to $w(ut)$ and $w(vt)$ but increasing with $i$ for $ut_{i}$ and decreasing with $i$ for $vt_{i}$. This way, if all $p^{2}$ copies of $t$ lie right of $u$ and $v$, we get a $p$-inversion among the $ut_{i}$’s or $vt_{i}$’s.
Figure 5: Illustration of the $2$-tree $H_{k}$ with edge-weighting $w_{k}$, for $k\geq 2$.
We then construct a sequence $\langle H_{1},w_{1}\rangle,\langle H_{2},w_{2}\rangle,\ldots$ of edge-weighted rooted $2$-trees, so that for every $k$, every left-growing vertex ordering of $\langle H_{k},w_{k}\rangle$ contains a $k$-inversion. The construction of $\langle H_{k},w_{k}\rangle$ is recursive, starting with $\langle H_{1},w_{1}\rangle$ being just a single edge of any weight. For $k\geq 2$, we start with $d=k^{2}$ vertices $a_{1},\ldots,a_{d}$ stacked onto the root edge $e^{*}=v^{\prime}v^{\prime\prime}$, and stacking a vertex $b^{\prime}_{i}$ onto each $v^{\prime}a_{i}$, as well as a vertex $b^{\prime\prime}_{i}$ onto $v^{\prime\prime}a_{i}$ for $i=1,\ldots,d$. All these edges have weight 00. Then, each $b^{\prime}_{i}a_{i}$ and $b^{\prime\prime}_{i}a_{i}$ is used as the base edge of two separate copies of $\langle H_{k-1},w_{k-1}\rangle$, where however the edge-weights $w_{k-1}$ of each copy are scaled and shifted to be in a specific half-open interval $[x,y)\subset\mathbb{R}$ depending on the current base edge ($b^{\prime}_{i}a_{i}$ or $b^{\prime\prime}_{i}a_{i}$). See Figure 5 for an illustration.
By induction, there is a $(k-1)$-inversion in each such copy of $H_{k-1}$. Using the pigeon-hole principle, and the Erdős–Szekeres theorem, we then identify an edge $a_{i}v^{\prime}$ (or $a_{i}v^{\prime\prime}$) which forms a $k$-inversion either together with $(k-1)$-inversion of one copy of $H_{k-1}$, or with $k-1$ edges, each from the $(k-1)$-inversion of a different copy of $H_{k-1}$.
To finish the proof, it is then enough to take $\langle H_{p},w_{p}\rangle$ as constructed above, and then apply 1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
As graphs of treewidth $2$ are always planar graphs, we conclude with the following result that contrasts similar results for the queue and stack number [22, 42]:
Corollary 5.5.
The priority queue number of planar graphs is unbounded. In particular, there is a planar graph $G$ with $n$ vertices and $\operatorname{pqn}(G)\geq\log_{4}n$.
Proof 5.6.
We observe that $H_{k}$ in the proof of Theorem 5.3 contains $4$ copies of $H_{k-1}$ while $H_{1}$ contains $2$ vertices. Thus, there are at least $4^{p}$ vertices in $G:=H_{p}$, which has treewidth $p$.
6 Complexity of Fixed-Ordering PQ-Layouts
[…]
Theorem 6.1.
[…]
Proof 6.2.
[…]
For a given circular-arc graph $H$, we next describe how to obtain an edge-weighted graph $\langle G,w\rangle$ and a vertex ordering $\sigma$ such that a $k$-coloring of $H$ directly corresponds to a PQ-layout of $\langle G,w\rangle$ under $\sigma$ with $k$ priority queues. Without loss of generality, we assume that we have a circular-arc representation of $H$ where all endpoints of the arcs are distinct.
We “cut” the circle at some point, which gives an interval representation where some vertices $S$ of $H$ refer to two intervals; in Figure 6, the circular arcs $a$, $b$, and $d$ are cut into intervals $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, and $d_{1}$ and $d_{2}$, respectively. We let the endpoints of all intervals be the vertices of $G$, whose ordering $\sigma$ along the spine is taken from the ordering of the corresponding endpoints of the intervals (for the intervals belonging to $S$, the order is arbitrary). Further, for each interval, we have an edge in $G$ connecting its two endpoints. To obtain $w$, we assign to the edges of $G$ the numbers 1 to $m$, where $m=|V(H)|+|S|$, in decreasing order of the right endpoints of the intervals. This way, every pair of edges whose corresponding intervals share a common point need to be assigned to distinct priority queues.
Figure 6: Illustration of our reduction.
So far, there is no mechanism that assures that $a_{1}$ and $a_{2}$, $b_{1}$ and $b_{2}$, etc., are assigned to the same priority queue. To this end, we add, for each such pair $\langle a_{1},a_{2}\rangle$, an edge from the left endpoint of $a_{1}$ to the right endpoint of $a_{2}$. We call these $|S|$ new edges synchronization edges, and we assign them the weights $m+1,\dots,m+|S|$ in decreasing order of their right endpoints. Furthermore, we add $k-1,k-2,\dots$ heavy edges below the leftmost, second leftmost, … vertex on the spine, respectively, as well as, $k-1,k-2,\dots$ heavy edges below the rightmost, second rightmost, … vertex on the spine, respectively, see Figure 6. Clearly, $k\geq|S|$, otherwise we have a no-instance. The heavy edges have distinct right endpoints and their weights are chosen such that (i) they are greater than $M$, where $M=m+|S|$ is the largest weight used so far, and (ii) they are chosen in decreasing order from left to right such that each pair of heavy edges needs to be assigned to distinct priority queues if they overlap.
[…]
Let $\Gamma$ be a PQ-layout of $\langle G,w\rangle$ under vertex ordering $\sigma$ having $k$ priority queues. For each vertex $a\in S$ ($\subseteq V(H)$), the two corresponding edges $a_{1}$ and $a_{2}$ in $G$ are assigned to the same priority queue in $\Gamma$.
{claimproof}
[Proof Sketch] Consider the vertices of $G$ from the outside to the inside. At the outer endpoint of $a_{1}$ ($a_{2}$, resp.), the edge $a_{1}$ ($a_{2}$, resp.) and the synchronization edge get the same color by the pigeonhole principle because all but one priority queue is occupied by the heavy edges and the previously considered edges, which all induce pairwise forbidden configurations.
If we have a PQ-layout of $\langle G,w\rangle$ under vertex ordering $\sigma$ with $k$ priority queues, we obtain, due to the choice of $w$ and $\sigma$, a coloring of $H$ by using the priority queues as colors. In particular, due to 2, both parts of the “cut” circular arcs get the same color. Conversely, we can easily assign the edges of $\langle G,w\rangle$ to $k$ priority queues if we are given a $k$-coloring of $H$. Clearly, our reduction can be implemented to run in polynomial time.
7 Open Problems
We conclude by summarizing a few intriguing problems that remain open. (i) Improve the upper and lower bounds for $\operatorname{pqn}(K_{n})$. (ii) Determine the complexity of deciding $\operatorname{pqn}(G)\leq k$ for $k>1$, or deciding $\operatorname{pqn}(G,w)\leq k$ for a given edge weighting $w$ when $k$ is a constant or no vertex ordering is given. (iii) Find graph families with bounded priority queue number; what about outerplanar graphs? (iv) Can the edge density of a graph $G$ be upper bounded by a term in $\operatorname{pqn}(G)$? (v) One can also investigate the structure of edge-weighted graphs $\langle G,w\rangle$ with a fixed vertex-ordering $\prec$. Edge-partitions of $G$ into $k$ priority queues are equivalent to proper $k$-colorings of an associated graph $H$, while $t$-inversions correspond to $t$-cliques in $H$. Is $\chi(H)$ bounded by a function in terms of $\omega(H)$? Is determining $\omega(H)$NP-complete?
Appendix
Appendix A Omitted Proofs of Section 3
See 3.3
Proof A.1.
Let $A$ and $B$ denote the two independent sets of size $n$. We partition the edge set into $n$ perfect matchings $M_{1},\ldots,M_{n}$. For each edge $e\in M_{i}$ we set $w(e)=i$. Thus, each vertex is incident to exactly one edge of weight $i$ for $i\in\{1,\ldots,n\}$.
Let $\Gamma$ be any PQ-layout of $K_{n,n}$. We first show that there exists a subgraph $K_{\frac{n}{2},\frac{n}{2}}$ with independent sets $A^{\prime}\subseteq A$ and $B^{\prime}\subseteq B$ of size $\frac{n}{2}$ such that the restriction $\Gamma^{\prime}$ of $\Gamma$ to the subgraph induced by $A^{\prime}\cup B^{\prime}$ is a separated PQ-layout. [^4] For simplicity, we assume that $n$ is even. Let $v_{1},\ldots,v_{t}$ be the first $t$ vertices in the linear ordering of $\Gamma$, such that |$|\{v_{1},\ldots,v_{t}\}\cap A|=\frac{n}{2}$ and $t$ is minimum. If |$|\{v_{t+1},\ldots,v_{2n}\}\cap B|\geq\frac{n}{2}$, we set $A^{\prime}=\{v_{1},\ldots,v_{t}\}\cap A$ and $B^{\prime}=\{v_{q},\ldots,v_{2n}\}\cap B$ where $q\geq t+1$ is chosen so that $|B^{\prime}|=\frac{n}{2}$. Otherwise, by the choice of $t$, we have that |$|\{v_{t+1},\ldots,v_{2n}\}\cap A|=\frac{n}{2}$, and, by the pigeon-hole principle, also |$|\{v_{1},\ldots,v_{t}\}\cap B|>\frac{n}{2}$. In this case, we choose $B^{\prime}=\{v_{1},\ldots,v_{p}\}\cap B$ where $p<t$ is chosen so that $|B^{\prime}|=\frac{n}{2}$ and $A^{\prime}=\{v_{t+1},\ldots,v_{2n}\}\cap A$. This proves the existence of $\Gamma^{\prime}$.
We now focus on $\Gamma^{\prime}$. Without loss of generality, we can assume that $A^{\prime}$ precedes $B^{\prime}$ in the linear ordering of $\Gamma^{\prime}$. Let $b_{1},\ldots,b_{\frac{n}{2}}$ denote the linear ordering of the vertices of $B^{\prime}$. Since $\Gamma^{\prime}$ is a separated PQ-layout and since each of its edges has an endpoint in $A^{\prime}$ and an endpoint in $B^{\prime}$, all edges of $\Gamma^{\prime}$ are co-active at $b_{1}$. As $K_{\frac{n}{2},\frac{n}{2}}$ is a subgraph of $K_{n,n}$, each vertex in $B^{\prime}$ is incident to exactly $\frac{n}{2}$ different edge weights occurring in the range $[1,n]$. Moreover, an edge $e$ incident to $b_{i}$ and an edge $e^{\prime}$ incident to $b_{j}$ form a forbidden configuration if $i<j$ but $w(e)>w(e^{\prime})$.
(a) The dark red cell represents the edge of weight 8, having vertex $b_{4}$ as its right endpoint. The red shaded cells represent the region $\mathcal{R}(e_{i})$.
(b) Let the black and red cells be set to 1. The red cells constitute a strictly monotonically decreasing path. Together, they represent a $k$-inversion in the corresponding edge-weighted graph.
(c) Let the black, red, and blue cells be set to 1. The red cells represent the candidates for edge $e_{1}$, while the blue edges represent the candidates for edge $e_{2}$.
Figure 7: Illustrations for the proof Theorem 3.3. In this matrix of size $n\times\frac{n}{2}$, all possible edge weights appear in the rows and the indices of the vertices $b_{1},\dots,b_{\frac{n}{2}}$ appear in the columns.
We now show that for some $k$ there is a $k$-inversion, that is, there is a sequence of edges $e_{1},\ldots,e_{k}$ such that for $1\leq i<j\leq k$, $w(i)>w(j)$ and $b_{\eta(i)}<b_{\eta(j)}$ where $\eta(i)$ is the index of the endpoint of edge $e_{i}$ in $B^{\prime}$. Another way to see this is by visualizing the relation between $\eta(i)$ and $w(e_{i})$ as a binary matrix [^5] A similar, although not identical, representation is used by Alam et al. [35] to prove the mixed page number of $K_{n,n}$. $M$ of size $n\times\frac{n}{2}$ where $M[w,\eta]=1$ if there is an edge $e_{i}$ in $\Gamma^{\prime}$ with $w(e_{i})=w$ and $\eta(i)=\eta$, and $M[w,\eta]=0$ otherwise; see Figure 7. In this view of the problem, for a given edge $e_{i}$ (see the red cell in Figure 7(a)), forbidden configurations are created with any edge whose cell (with value $1$) occurs in the region $\mathcal{R}(e_{i})$ to the top-left of the cell of $e_{i}$ (see the red shaded cells in Figure 7(a)). Moreover, each column of the matrix has exactly $\frac{n}{2}$ cells with value $1$ (corresponding to the $\frac{n}{2}$ edges incident to each vertex in $B^{\prime}$). Our desired sequence $e_{1},\ldots,e_{k}$ thus is a strictly monotonically decreasing path through the matrix along cells with value $1$; see the red path in Figure 7(b). It is easy to see that each pair of edges along the sequence creates a forbidden configuration, i.e., $k$ priority queues are necessary.
To compute a strictly monotonically decreasing path of maximum length, we proceed iteratively as follows. At step 1 we compute a subset of candidate edges (cells) for $e_{1}$, containing to every cell of an edge $e$ for which $\mathcal{R}(e)$ is empty (i.e., it does not contain cells of value 1). At step $i$, with $i\geq 2$, we compute a subset of candidate edges (cells) for $e_{i}$, containing every cells of an edge $e$ for which $\mathcal{R}(e)$ is empty after all cells of the candidate edges for $e_{j}$ with $j<i$ have been removed (i.e., are set to 0 in the matrix). For example, in Figure 7(c), the red cells correspond to the candidate edges for $e_{1}$, while the blue cells correspond to the candidate edges for $e_{2}$. We iteratively repeat the process until all $\frac{n^{2}}{4}$ cells with value $1$ have been set to 0. The number $k$ of steps in this process yields the length of the sequence and hence the required number of priority queues.
We now compute the value of $k$. To this end, we first prove that, in step $i$, the topmost $i-1$ rows and the leftmost $i-1$ columns are empty (i.e., all cells are set to 0). The proof is by induction on $i$. When $i=1$ the claim is trivially true. Assume that the claim is true for $i-1$, with $i\geq 2$. For each edge $e$ whose cell is at column $i$, the region $\mathcal{R}(e)$ contains only cells at columns that are to the left of column $i$, which are empty by the inductive hypothesis. Hence, all cells at column $i$ are set to 0 in step $i$. An analogous argument applies for the topmost $i-1$ rows. Moreover, denote by $M^{\prime}$ the sub-matrix of $M$ obtained from $m$ after removing the $i-1$ leftmost columns and the $i-1$ topmost rows. In each diagonal of $M^{\prime}$ of slope $-1$, at most one non-empty cell has an empty region to the top-left. Observe that $M^{\prime}$ has $n-i+1$ rows and $\frac{n}{2}-i+1$ columns; thus the number of such diagonals is $\frac{3n}{2}-2i+1$ and the number of cells set to 0 at step $i$ is at most $\frac{3n}{2}-2i+1$. The number of iterations $k$ is the minimum integer that satisfies the following inequality:
$\sum_{i=1}^{k}\left(\frac{3n}{2}-2i+1\right)\geq\frac{n^{2}}{4}.$
We have:
$\sum_{i=1}^{k}\left(\frac{3n}{2}-2i+1\right)=\left(\frac{3n}{2}+1\right)k-2\sum_{i=1}^{k}i=\left(\frac{3n}{2}+1\right)k-k^{2}-k=\frac{3n}{2}k-k^{2}$
and, hence,
$k^{2}-\frac{3n}{2}k+\frac{n^{2}}{4}\leq 0.$
Equality is reached for
$k_{1,2}=\frac{3n}{4}\pm\sqrt{\frac{9n^{2}}{16}-\frac{4n^{2}}{16}}=\frac{3\pm\sqrt{5}}{4}n.$
This means that the inequality is satisfied for every value in the interval $I=[k_{1},k_{2}]$. Since $k$ must be the minimum integer value in $I$, we have $k=\left\lceil\frac{3-\sqrt{5}}{4}n\right\rceil$.
Appendix B Omitted Proofs of Section 4
See 4.1
Proof B.1.
To prove $\operatorname{pqn}(G^{\prime})=1$, we show that given any edge-weight function $w^{\prime}$ for the edges of $G^{\prime}$, there is a PQ-layout of $G^{\prime}$ with one priority queue. From $\langle G^{\prime},w^{\prime}\rangle$, we construct the edge-weighted graph $\langle G,w\rangle$ where $w$ is an edge-weight function for $V$ that is identical [^6] For edges incident to $u$ and $v$, we use the edge weights of the edges incident to $x_{uv}$. So, if there are two edges in $G$ that are represented by the same edge in $G^{\prime}$, they get the same weight. to $w^{\prime}$ except for having a weight of $W$ for edge $e_{0}$ such that $W$ is smaller than any other edge weight in $w$. Since $\operatorname{pqn}(G)=1$, there is a PQ-layout $\Gamma$ with one priority queue for $\langle G,w\rangle$. We next adjust $\Gamma$ slightly to obtain a PQ-layout $\Gamma^{\prime}$ with one priority queue for $\langle G^{\prime},w^{\prime}\rangle$. For an illustration of this process, see Figure 8.
(a) PQ-layout $\Gamma$ before contracting the edge $uv$.
(b) PQ-layout $\Gamma^{\prime}$ after contracting the edge $uv$.
Figure 8: Example of the procedure for maintaining a PQ-layout with one priority queue after edge contraction as described in the proof of Lemma 4.1.
W.l.o.g. let $u$ precede $v$ along the spine of $\Gamma$. Further, let $L$ be the ordered set of vertices to the left of $u$, let $M$ be the ordered set of vertices between $u$ and $v$, and let $R$ be the ordered set of vertices to the right of $v$. In $\Gamma^{\prime}$, we keep $L$ and $R$ unchanged and we place $M\cup\{x_{uv}\}$ in between as follows. First place $x_{uv}$, then add to the right each vertex $m\in M\cap N(x_{uv})$ (where $N(x_{uv})$ are the neighbors of $x_{uv}$) in increasing order with respect to the weight of its incident edge $x_{uv}m$, and finally add the remaining vertices of $M$ in arbitrary order.
It remains to prove that there is no forbidden configuration in $\Gamma^{\prime}$. No pair of edges having both their right endpoints in $L\cup R$ is involved in a forbidden configuration because otherwise, it would have been involved in a forbidden configuration in $\Gamma$ already. The crucial observation is that, in $\Gamma$, no vertex $m\in M$ is the right endpoint of some edge $e$ because the weight $w(e)$ would be greater than $W$, but $e$ would end below $e_{0}$, which has weight $W$. Therefore, the edges incident to vertices in $M$ have their neighbors only in $R\cup\{v\}$. After merging $u$ and $v$ to $x_{uv}$ and placing $x_{uv}$ to the left of all vertices in $M$, the vertices in $M\cap N(x_{uv})$ have exactly one incoming edge from the left, while the other vertices in $M$ have none. By sorting $M\cap N(x_{uv})$ increasingly by the weights of these incoming edges, we avoid a forbidden configuration. Moreover, no edge having its right endpoint in $M\cup\{x_{uv}\}$ is involved in a forbidden configuration with an edge having its right endpoint in $L\cup R$ because this would have been a forbidden configuration in $\Gamma$ already.
See 4.10
Proof B.2.
For a leaf $\ell$, let $p(\ell)$ denote its neighbor in $C(L)$ and let $e^{\star}$ be an edge of $C(L)$ such that $w(e^{\star})$ is maximum among all edges of $C(L)$. We remove $e^{\star}$ and every leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$ from $L$ obtaining a caterpillar $T$ where both degree-$1$ vertices of the underlying path $P(T)$ are endpoints of $e^{\star}$. We then find a PQ-layout of $T$ using Lemma 4.6 chosing one of the endpoints of $e^{\star}$ as $r$. Note that for each edge $e$ in $T$ we have that $w(e^{\star})\geq w(e)$. While $e^{\star}$ may be co-active with edges in $e$, we have that its right endpoint succeeds any vertex in $T$, i.e., no forbidden configuration is created when we reinsert $e^{\star}$. It remains to reinsert each leaf $\ell$ with $w({\ell}p(\ell))>w(e^{\star})$. Let $S$ denote the set of such leaves and for $\ell\in S$ let $w(\ell)=w({\ell}p(\ell))$. We insert all vertices of $S$ to the right of $r$ sorted in increasing order of weight. It remains to discuss that for $\ell\in S$, edge ${\ell}p(\ell)$ is not involved in forbidden configurations. Let $uv$ with $u\prec v$ be an co-active edge of ${\ell}p(\ell)$. If $v\not\in S$, we have that $v\prec\ell$ and $w(uv)\leq w(e^{\star})<w({\ell}p(\ell))$, i.e., no forbidden configuration occurs. Otherwise, $v$ and $\ell$ are sorted by their weights in $\prec$ which are equal to $w(uv)$ and $w({\ell}p(\ell))$, respectively. Again no forbidden configuration occurs and the statement follows.
Finding and removing a maximum-weight edge of the cycle and all edges with greater weight can be done in $\mathcal{O}(n)$ time. Arranging the resulting graph using Lemma 4.6 as well. For the subsequent reinsertion of edges, we need to sort the leaves by weight. This requires and can be done in $\mathcal{O}(n\log n)$ time, which is also the resulting running time of the algorithm.
See 4.14
Proof B.3.
Assume without loss of generality that $w(ad)\geq\max\{w(ab),w(bc),w(cd)\}$. Consider the edge $e_{c}$ incident to $c$ on $P(C_{c})$. Assume first, that $w(e_{c})\leq w(ad)$. For $\Box$, we use the layout $d\prec c\prec b\prec a$. Note that only the edge $ad$ nests and pseudo-nests edges of $\Box$, however it is the heaviest edge in $\Box$. Further, we use Lemma 4.4 for laying out $C_{a}$ with leftmost vertex $a$. The resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Now, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma 4.6 and place all vertices of $C_{c}^{\prime}$ except for $c$ before vertex $d$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ad$. We have that $c\prec a$ and $w(e_{c})\leq w(ad)$, thus this creates no forbidden configuration.
Next, assume that $w(e_{c})>w(ad)$. For $\Box$, we use the layout $a\prec b\prec c\prec d$. Again, only edge $ad$ nests and pseudo-nests edges of $\Box$, but still it is the heaviest edge in $\Box$. Further, we use Lemma 4.6 for laying out $C_{a}$ with rightmost vertex $a$. Once more, the resulting PQ-layout of $G$ without $C_{c}$ clearly has priority queue number 1. Again, consider the caterpillar $C_{c}^{\prime}$ obtained from $C_{c}$ by removing the leaves of $c$. We lay out $C_{c}^{\prime}$ using Lemma 4.4 and place all vertices of $C_{c}^{\prime}$ except for $c$ after vertex $d$ in $\prec$. Thus, there is one intersection between $C_{c}^{\prime}$ and $\Box$, namely, between $e_{c}$ and $ad$, and one pseudo-nesting, namely between $e_{c}$ and $cd$. However, we have that $d\prec c$ and $w(e_{c})>w(ad)\geq w(cd)$, thus this creates no forbidden configuration.
Finally, in both cases, we only have to reintroduce the leaves $L_{c}$ of vertex $c$ in $C_{c}$ to complete the layout. For $\ell\in L_{c}$, let $w(\ell)=w({\ell}c)$. We iteratively insert the leaves of $c$ with increasing weight. If $w(\ell)\leq w(ad)$, we place $\ell$ directly before $c$ in $\prec$. Then, edge ${\ell}c$ only creates a nesting with edge $ad$, however $ad$ has heavier weight, so no forbidden configuration is introduced. Otherwise, $w(\ell)>w(ad)$ and we find a position for $\ell$ as follows. Let $C_{r}$ be the caterpillar whose non-root vertices are drawn to the right of the vertices of $\Box$, i.e., in the first case $C_{a}$ and in the second case $C_{c}^{\prime}$. Also let $v^{\star}$ be the leftmost vertex of the caterpillar $C_{r}$ with parent $p(v^{\star})$ in $C_{r}$ for which $w(v^{\star}p(v^{\star}))>w(\ell)$. Then, we place $\ell$ directly before $v^{\star}$ in $\prec$. For a contradiction, assume that this creates a forbidden configuration which clearly must involve edge ${\ell}c$ and another edge $uv$ with $u\prec v$ as our procedure ensures that the edges between $c$ and its leaves which pseudo-nest are not creating forbidden configurations. Observe that $uv$ is not an edge of $\Box$ as all edge weights of $\Box$ are smaller than the weight of edge ${\ell}c$ and $\ell$ is placed to the right of all vertices of $\Box$. Thus $u=p(v)$ and if $v\prec\ell$ and $w({\ell}c)<w(uv)$, we have a contradiction to our choice for the position of $\ell$. Thus, we must have $\ell\prec v$ and $w({\ell}c)>w(uv)$. By our choice of the position of $\ell$, we have that $v^{\star}\prec v$. Since also $u\prec\ell$, it follows that $uv$ and $p(v^{\star})v^{\star}$ are co-active, however, this is a contradiction to the fact that $C_{r}$ is drawn with a priority-queue-number-one layout. The proof follows.
See 4.17
(a) Case 1
(b) Case 2.1
(c) Case 2.2
(d) Case 3
Figure 9: Illustration for the proof of Lemma 4.17.
Proof B.4.
Let $G=K_{2,3}$ and consult Figure 9 for an illustration. Let $\{u_{1},u_{2}\}$ be the vertex partition with vertices of degree 3 and let $\{v_{1},v_{2},v_{3}\}$ be the vertex partition with vertices of degree 2. Let $w$ be any given edge-weight function and denote by $e_{1}$ the heaviest edge of $\langle G,w\rangle$ and by $e_{2}$ the second heaviest edge. We distinguish three main cases and, for each of them, we describe how to construct a PQ-layout $\Gamma$ such that $\operatorname{pqn}(\Gamma)=1$.
  • Case 1: $e_{1}$ and $e_{2}$ are disjoint. Refer to Figure 9(a). We may assume that $e_{1}=u_{1}v_{1}$ and $e_{2}=u_{2}v_{3}$, and define the vertex ordering as $v_{3}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{1}$. Since $e_{1}$ and $e_{2}$ are the two heaviest edges, there are neither forbidden nestings nor forbidden pseudo-nestings in $\Gamma$. Also, the only two crossing edges are $e_{1}$ and $e_{2}$, and $w(e_{2})\leq w(e_{1})$. Since $e_{2}$ is pulled from the priority queue before $e_{1}$, there is no forbidden crossing configuration.
  • Case 2: $e_{1}$ and $e_{2}$ share a degree-2 vertex. Let $e_{3}$ be the third heaviest edge of $\langle G,w\rangle$. We consider two further subcases:
    • Case 2.1: $e_{3}$ shares a vertex with $e_{1}$. Refer to Figure 9(b). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{1}v_{1}$. The linear ordering of $\Gamma$ is set as $v_{1}\prec u_{2}\prec v_{2}\prec u_{1}\prec v_{3}$. As for Case 1, there are no forbidden (pseudo-)nesting configurations. Also, the only two crossing edges are $e_{2}$ and $e_{3}$, and $w(e_{3})\leq w(e_{2})$. Since $e_{3}$ is pulled from the queue before $e_{2}$, there is no forbidden crossing configuration.
    • Case 2.2: $e_{3}$ shares a vertex with $e_{2}$. Refer to Figure 9(c). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$, $e_{2}=u_{2}v_{3}$, and $e_{3}=u_{2}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{1}\prec u_{1}\prec v_{2}\prec u_{2}\prec v_{3}$. The argument is similar to the previous case.
  • Case 3: $e_{1}$ and $e_{2}$ share a degree-3 vertex. Refer to Figure 9(d). Assume, w.l.o.g., that $e_{1}=u_{1}v_{3}$ and $e_{2}=u_{1}v_{1}$. We define the linear ordering of $\Gamma$ as $v_{3}\prec v_{1}\prec u_{2}\prec v_{2}\prec u_{1}$. As in the previous cases, for the choice of $e_{1}$ and $e_{2}$, there are no forbidden nesting or pseudo-nesting configurations. Also, the only two crossing edges are $e_{2}$ and the edge $v_{3}u_{2}$; since $v_{3}u_{2}$ is pulled before $e_{2}$ from the priority queue, and since $w(v_{3}u_{2})\leq w(e_{2})$, there is no forbidden crossing configuration.
See 4.19
Proof B.5.
Suppose for a contradiction that there is a PQ-layout $\Gamma$ of $\langle C,w\rangle$ with $\operatorname{pqn}(\Gamma)=1$ where the last vertex is incident to two edges that both do not have the maximum weight. Let $e=uv$ be an edge of maximum weight with $u\prec v$. (If there are multiple edges with maximum weight, pick $e$ such that $v$ is rightmost among all right endpoints of those heaviest edges.) We claim that there is an edge $e^{\prime}=xy$ that is co-active with $e$ and $v\prec y$ resulting in a forbidden configuration since $w(e)>w(e^{\prime})$. Let $a$ be the leftmost vertex and $z$ be rightmost vertex of $\Gamma$. There are two vertex-disjoint paths between $a$ and $z$. Only one of these paths can contain $v$, while the other needs to bridge $v$ – this implies that $e^{\prime}$ exists.
See 4.20
Proof B.6 (Proof of Cases $\mathcal{F}_{4}$$\mathcal{F}_{8}$).
It remains to analyze $\mathcal{F}_{4}$$\mathcal{F}_{8}$. For an illustration, refer to Figure 4.
$\mathcal{F}_{4}$, $\mathcal{F}_{5}$
: By Lemma 4.19 and symmetry, we may assume that $a,b\prec c$. Since each edge not in $\triangle abc$ has a weight strictly between $w(ab)=w(ac)$ and $w(bc)$, no such edge can start weakly left of $\min\{a,b\}$ and end strictly right of $\min\{a,b\}$. In particular, it follows that $d\prec a$ and, since also $w(ed)<w(ad)$, that $e\prec d$. In summary, we have $e\prec d\prec a\prec c$ and $b\prec c$.
Now for $\mathcal{F}_{4}$ it follows that the right endpoint of edge $fg$ must be left of $e$. But then edge $ed$ nests below $af$; a contradiction, as $w(ed)>w(af)$.
For $\mathcal{F}_{5}$, first note that if $a\prec b$, then either edge $bf$ ends below $ac$ (if $f\prec c$) or edge $bc$ ends below $bf$ (if $c\prec f$); either is not allowed. So we have $b\prec a$, and hence $f\prec a$, as argued in the beginning. But now either edge $bf$ ends below $ad$ (if $d\prec b$) or edge $ed$ ends below $ab$ (if $b\prec d$); either is not allowed.
$\mathcal{F}_{6}$, $\mathcal{F}_{7}$
: By Lemma 4.19, $b$ or $c$ is the last vertex of the 4-cycle $abcd$ and $a$ or $c$ is the last vertex of $\triangle abc$. Hence, $a,b,d\prec c$. Moreover, $a\prec d$ since otherwise, $ad$ would end below the lighter edge $cd$. We know that $e\prec c$ because otherwise $ac$ would end below the lighter edge $de$ ($ae$, respectively).
For $\mathcal{F}_{6}$, further $e\prec d$ because if $d\prec e\prec c$, then $de$ would end below the lighter edge $cd$. It remains to place $b$ (we already know $b\prec c$). It cannot be $b\prec d$ because then $de$ would end below the lighter edge $bc$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
For $\mathcal{F}_{7}$, a similar argument applies. We have $e\prec a$ because if $a\prec e\prec c$, then $ae$ would end below one of the lighter edges $ad$ or $cd$. If $b\prec a\prec d\prec c$, then $ae$ would end below the lighter edge $bc$. If $a\prec b\prec d\prec c$, then $ab$ would end below the lighter edge $ad$. Hence, we have $a\prec d\prec b\prec c$. Now, however, $ab$ ends below the lighter edge $cd$; a contradiction.
$\mathcal{F}_{8}$
: By Lemma 4.19, the last vertex of the 4-cycle $abcd$ is $a$ or $b$. In the former case, also $c\prec d$, as otherwise $cd$ ends below the lighter edge $ad$, and $b\prec d$, as otherwise $cd$ ends below the lighter edge $bc$. This leaves four possible orderings: $b\prec c\prec d\prec a$, $c\prec b\prec d\prec a$, and, symmetrically if $b$ is the last vertex, $a\prec d\prec c\prec b$, and $d\prec a\prec c\prec b$.
First suppose that $b\prec c\prec d\prec a$. Then $c\prec f\prec a$ because $df$ has the lowest weight and there are heavier edges ending at $c$ and $a$. Now $f$ cannot be the right endpoint of $fg$ because then the edge $fg$ would end below one of the lighter edges $cd$ or $ad$. However, $f$ can also not be the left endpoint of $fg$ because then the edge $fg$ would end below the lighter edges $cd$ or $ad$, or the heavier edge $ab$ would end below $fg$.
Similarly, if $c\prec b\prec d\prec a$, then $b\prec f\prec a$ by the previous argument. Again, $f$ cannot be the right endpoint of $fg$ because then $fg$ would end below $cd$ or $ad$, and it cannot be the left endpoint of $fg$ by the same argument as before.
Now suppose that $a\prec d\prec c\prec b$. By a similar argument as before, having $a\prec f\prec c$ leads to at least one forbidden configuration. Here, however, $f\prec a$ is possible because there is no edge ending at $a$ that could form a forbidden configuration with the lightest edge $df$ It follows that $a\prec e$ and $g\prec f$. It remains to give $e$ a precise position. If $a\prec e\prec b$, then $ae$ ends below one of the lighter edges $ad$, $cd$, or $bc$. However, if $b\prec e$, then the heavier edge $ab$ ends below $ae$.
Finally suppose that $d\prec a\prec c\prec b$. We can argue again that $g\prec f\prec d\prec a\prec c\prec b$. It remains to place $e$. Clearly, $a\prec e$ as otherwise $ae$ ends below $cd$. Now, however, $ae$ ends below the lighter edges $cd$ or $cb$, or the heavier edge $ab$ ends below $ae$; a contradiction.
See 4.22
Proof B.7.
Let $G=(V,E)$ be any fixed graph. Our task is to show that either $\operatorname{pqn}(G)=1$, or $G$ contains some $\mathcal{F}\in\{\mathcal{F}_{1},\ldots,\mathcal{F}_{8}\}$ as a minor. For this, we may assume that $G$ is connected. If $G$ is a tree, i.e., contains no cycles, then $\operatorname{pqn}(G)=1$ by Lemma 4.4 and we are done.
Next assume that $G$ contains exactly one cycle $C$. Then $G-E(C)$ is a forest $F$, and we consider each component of $F$ as rooted at the vertex of $C$. For a component $K$ of $F=G-E(C)$ we say that $K$ is
  • trivial if $K$ is just an isolated vertex,
  • a rooted star if $K$ is a star whose central vertex lies on $C$,
  • a rooted caterpillar if $K$ is a caterpillar with an underlying path with one endpoint on $C$,
  • complex if none of the above applies for $K$.
First assume that some component $K$ of $F$ is complex. Let $w$ be the vertex in $K$ that lies on $C$. If $K$ is not a caterpillar, then $K$ contains a vertex $v$ (possibly $w$) with at least three neighbors of degree at least $2$ each. Contracting $C$ to a triangle and the $v$-$w$-path in $K$ to a single vertex results in a graph that contains $\mathcal{F}_{4}$. Hence, $G$ contains $\mathcal{F}_{4}$ as a minor. So assume that $K$ is a caterpillar but no underlying path of $K$ starts with $w$. Let $P$ be the path in $K$ on all non-leaf vertices of $K$. Then $w$ is neither an endpoint of $P$ nor adjacent to an endpoint of $P$. It follows that either $w$ is an interior vertex of $P$ or $w$ is adjacent to an interior vertex $w^{\prime}$ of $P$. Contracting $C$ again to a triangle, and contracting the edge $ww^{\prime}$ in the latter case, we again obtain a graph that contains $\mathcal{F}_{4}$. Hence again, $G$ contains $\mathcal{F}_{4}$ as a minor and we are left with the case that no component of $F$ is complex.
Now if every component of $F$ is trivial or a rooted star, then $\operatorname{pqn}(G)=1$ by Lemma 4.10. So assume some component $K$ of $F$ is a rooted caterpillar. If all other components (different from $K$) are trivial, then $\operatorname{pqn}(G)=1$ by Lemma 4.12. If $F$ has besides $K$ two further non-trivial components $K^{\prime},K^{\prime\prime}\neq K$, then $G$ contains $\mathcal{F}_{5}$ as a minor. So assume now that there is exactly one further non-trivial component $K^{\prime}\neq K$. If the roots of $K$ and $K^{\prime}$ have distance at least $3$ in one direction along $C$, then $G$ contains $\mathcal{F}_{8}$ as a minor. So, if none of the above is the case, then either $C$ is a triangle and $\operatorname{pqn}(G)=1$ by Corollary 4.16, or $C$ is a quadrangle and the roots of $K$ and $K^{\prime}$ are opposite on $C$, and $\operatorname{pqn}(G)=1$ by Lemma 4.14.
The case where $G$ contains at least two cycles remains. First assume that $G$ contains two edge-disjoint cycles $C_{1},C_{2}$. If $C_{1},C_{2}$ have at most one vertex in common, then $G$ contains $\mathcal{F}_{2}$ as a minor. If $C_{1},C_{2}$ have at least two vertices in common, we claim that $G$ contains $\mathcal{F}_{7}$ as a minor. One of $C_{1},C_{2}$, say $C_{2}$, has length at least $4$. Let $a$ and $c$ be two common vertices of $C_{1},C_{2}$ (consider Figure 4(g) for the naming scheme). Then $C_{1}$ can be contracted to the triangle $abc$, while $C_{2}$ and can be contracted to the 4-cycle $adce$. Omitting edge $ce$ yields $\mathcal{F}_{7}$.
Finally, consider the case that any two cycles in $G$ share at least one edge. In fact, either $G$ contains $\mathcal{F}_{1}=K_{4}$ as a minor (and we are done) or there are two vertices $x,y$ in $G$ that are both contained in all cycles. Since there are no edge-disjoint cycles, there are at most three edge-disjoint $x$-$y$-paths in $G$. As $G$ has two cycles, there are exactly three such $x$-$y$-paths, say $P_{1},P_{2},P_{3}$. If $G$ is not just $P_{1}\cup P_{2}\cup P_{3}$, then $G$ contains $\mathcal{F}_{6}$ or $\mathcal{F}_{7}$ as a minor. If at least one of $P_{1},P_{2},P_{3}$ has at least three edges, then $G$ contains $\mathcal{F}_{3}$ as a minor. So, if none of the above is the case, then either $G=K_{2,3}$ or $G=K_{4}-e$ and we have $\operatorname{pqn}(G)=1$ by Lemmas 4.17 and 4.18.
Appendix C Omitted Content of Section 5
Here, we provide a formal definition of treewidth and pathwidth. Given an undirected graph $G=(V,E)$, a tree decomposition is a tree $T$ with nodes $X_{1},\dots,X_{k}$ such that:
  • for each $i\in\{1,\dots,k\}$, $X_{i}\subseteq V$,
  • $\bigcup_{i\in\{1,\dots,k\}}X_{i}=V$,
  • if $X_{i}$ and $X_{j}$ both contain a vertex $v\in V$, then all nodes of $T$ that lie on the (unique) path between $X_{i}$ and $X_{j}$ contain $v$ as well, and
  • for every $uv\in E$, there is some $i\in\{1,\dots,k\}$ such that $\{u,v\}\subseteq X_{i}$.
The width of a tree decomposition is $\max\{|X_{i}|-1\ :i\in\{1,\dots,k\}\}$, i.e., it is the size of its largest set $X_{i}$ minus one. The treewidth of $G$ is the minimum width among all possible tree decompositions of $G$. The pathwidth is defined similarly. The only difference is that $T$ needs to be a path, which is hence known as path decomposition.
Now, we present the full proof of Theorem 5.3.
See 5.3
Proof C.1.
Let $p\geq 1$ be a fixed integer. Our desired graph $G_{p}$ will be a rooted $2$-tree, i.e., an edge-maximal graph of treewidth $2$ with designated root edge, which are defined inductively through the following construction sequence:
  • A single edge $e^{*}=v^{\prime}v^{\prime\prime}$ is a $2$-tree rooted at $e^{*}$.
  • If $G$ is a $2$-tree rooted at $e^{*}$ and $e=uv$ is any edge of $G$, then adding a new vertex $t$ with neighbors $u$ and $v$ is again a $2$-tree rooted at $e^{*}$. In this case we say that $t$ is stacked onto the edge $uv$, and vertices $u$ and $v$ are called the parents of $t$.
We define a vertex ordering $\sigma$ of a $2$-tree rooted at $e^{*}=v^{\prime}v^{\prime\prime}$ to be left-growing if no vertex (different from $v^{\prime}$, $v^{\prime\prime}$) appears to the right of both its parents.
See 1
{claimproof}
Given $\langle G,w\rangle$, let us fix some $\varepsilon>0$ smaller than the minimum difference between any two distinct edge-weights in $\langle G,w\rangle$, i.e., $0<\varepsilon<\min\{|w(e_{1})-w(e_{2})|\colon e_{1},e_{2}\in E(G),w(e_{1})\neq w(e_{2})\}$. Let us call a weight function $w^{\prime}\colon E(G)\to\mathbb{R}$ to be $\varepsilon$-close to $w$ if $|w^{\prime}(e)-w(e)|\leq\varepsilon/2$ for every edge $e\in E(G)$. Clearly, for any fixed vertex ordering, a $p$-inversion for edge weights $w$ is also a $p$-inversion for $w^{\prime}$, if $w^{\prime}$ is $\varepsilon$-close to $w$.
Now consider the given $\langle G,w\rangle$ and a construction sequence of $G$ starting with the root edge $e^{*}$. We shall construct $2$-tree $\tilde{G}$ such that for each edge $e$ in $G$ there is a corresponding set of edges in $\tilde{G}$, which we call copies of $e$. It starts with exactly $1=p^{0}=p^{2\cdot 0}$ copies of the root edge $e^{*}$. Then, whenever a vertex $t$ is stacked onto an edge $e=uv$ in the construction sequence of $G$, for each existing copy $\tilde{e}=\tilde{u}\tilde{v}$ of $e$ we stack $p^{2}$ vertices $t_{1},\ldots,t_{p^{2}}$ on $\tilde{e}$ in $\tilde{G}$. As edge-weights we choose
$\tilde{w}(\tilde{u}t_{i})=w(ut)+i\cdot\frac{\varepsilon}{2p^{2}}\qquad\text{and}\qquad\tilde{w}(\tilde{v}t_{i})=w(vt)-i\cdot\frac{\varepsilon}{2p^{2}}\qquad\text{for $i=1,\ldots,p^{2}$.}$
Each such edge $\tilde{u}t_{i}$ is a copy of $ut$, while each such $\tilde{v}t_{i}$ is a copy of $vt$. Thus, if there are $p^{2j}$ copies of $e$ in $\tilde{G}$, then there are $p^{2(j+1)}$ copies of $ut$ and $p^{2(j+1)}$ copies of $vt$ in $\tilde{G}$. This completes the construction of $\langle\tilde{G},\tilde{w}\rangle$.
Now fix any vertex ordering $\sigma$ of $\tilde{G}$. Consider a vertex $t$ that is stacked onto an edge $e=uv$ of $G$, and one copy $\tilde{e}=\tilde{u}\tilde{v}$ of $e$ in $\tilde{G}$. If all $p^{2}$ corresponding vertices $t_{1},\ldots,t_{p^{2}}$ in $\tilde{G}$ lie right of both vertices of $\tilde{e}$ in $\sigma$, then by the Erdős–Szekeres theorem there is a subset of at least $p$ such vertices with all decreasing or all increasing indices in the left-to-right order in $\sigma$. In the former case the corresponding copies of $ut$ form a $p$-inversion, in the latter case the corresponding copies of $vt$ form a $p$-inversion.
Thus we may assume that for any copy of any edge $e$ of $G$ and any vertex $t$ stacked on $e$, at least one corresponding vertex $t_{i}$ does not appear to the right of both of its parents in $\tilde{G}$. But then there is a subgraph in $\tilde{G}$ isomorphic to $G$ for which $\sigma$ is left-growing and whose restriction $w^{\prime}$ of $\tilde{w}$ is $\varepsilon$-close to $w$. By assumption, this subgraph contains a $p$-inversion.
Now we construct a sequence $\langle H_{1},w_{1}\rangle,\langle H_{2},w_{2}\rangle,\ldots$ of edge-weighted rooted $2$-trees, so that for every $k$, every left-growing vertex ordering of $\langle H_{k},w_{k}\rangle$ contains a $k$-inversion.
Claim 3.
For each $k\geq 1$, any given edge $e^{*}$ of any given weight $w^{*}$, and any non-empty half-open interval $[x,y)\subset\mathbb{R}$, there exists a $2$-tree $H_{k}$ rooted at $e^{*}$ and an edge weighting $w_{k}$ of $H_{k}$ such that
  • $w_{k}(e^{*})=w^{*}$ while $w_{k}(e)\in[x,y)$ for each $e\in E(H_{k})-e^{*}$,
and in every left-growing vertex ordering $\sigma$ of $\langle H_{k},w_{k}\rangle$ there is
  • a $k$-inversion $(e_{1},\ldots,e_{k})$ without $e^{*}$, but whose right end is the right endpoint of $e^{*}$.
{claimproof}
We proceed by induction on $k$. For $k=1$, it is enough to let $H_{k}$ be just the given root edge $e^{*}$ (with $w_{k}(e^{*})=w^{*}$) and one vertex $t$ stacked onto $e^{*}$ with both edge-weights in $[x,y)$. In any left-growing vertex ordering, the edge between $t$ and the right endpoint of $e^{*}$ is the desired $1$-inversion.
So for the remainder let us fix $k\geq 2$, and assume that the claim holds for $k-1$. Our task is to construct $\langle H_{k},w_{k}\rangle$ for the given root edge $e^{*}=v^{\prime}v^{\prime\prime}$, given weight $w^{*}$, and given interval $[x,y)$. By additive and multiplicative shifts of the weight function, we can assume for convenience that $[x,y)=[0,d+1)$ for $d=2k^{2}$. Starting with $e^{*}=v^{\prime}v^{\prime\prime}$ of weight $w_{k}(e^{*})=w^{*}$, we construct $\langle H_{k},w_{k}\rangle$ as follows. See Figure 5 for an illustration.
  • Stack $d$ vertices $a_{1},\ldots,a_{d}$ onto the root edge $e^{*}$, giving weight 00 to all $2d$ new edges.
  • For each $i=1,\ldots,d$
    • stack a vertex $b^{\prime}_{i}$ onto edge $a_{i}v^{\prime}$ giving weight 00 to both new edges,
    • let $a_{i}b^{\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[d+1-i,d+1-i+1)$,
    • stack a vertex $b^{\prime\prime}_{i}$ onto edge $a_{i}v^{\prime\prime}$ giving weight 00 to both new edges, and
    • let $a_{i}b^{\prime\prime}_{i}$ be the root edge of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[i,i+1)$, as well as the root of $\langle H_{k-1},w_{k-1}\rangle$ with interval $[d+1-i,d+1-i+1)$.
Now consider any left-growing vertex ordering $\sigma$ of $H_{k}$. Since the construction is symmetric, we may assume without loss of generality that $v^{\prime}$ is the right endpoint of the root edge $e^{*}$ in $\sigma$. Hence, each of $a_{1},\ldots,a_{d},b^{\prime}_{1},\ldots,b^{\prime}_{d}$ lies left of $v^{\prime}$ since $\sigma$ is left-growing and $v^{\prime}$ is a parent of all these vertices. By the pigeon-hole principle, there is a set $I_{1}\subseteq[d]$ of at least $d/2=k^{2}$ indices such that either $a_{i}\prec b^{\prime}_{i}$ for all $i\in I_{1}$ or $b^{\prime}_{i}\prec a_{i}$ for all $i\in I_{1}$. By symmetry [^7] We are not using in the remainder that $a_{i}v^{\prime\prime}$ is an edge, while $b^{\prime}_{i}v^{\prime\prime}$ is not., let us assume that $b^{\prime}_{i}\prec a_{i}$ for all $i\in I_{1}$.
By the Erdős–Szekeres theorem, there is a subset $I_{2}\subseteq I_{1}$ of at least $\sqrt{k^{2}}=k$ indices such that either $a_{i}\prec a_{j}$ for all $i,j\in I_{2}$ with $i<j$ or $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with $i<j$. By symmetry, let us assume that $a_{j}\prec a_{i}$ for all $i,j\in I_{2}$ with $i<j$ [^8] We use the symmetry of the two copies of $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ here. In the assumed case, we focus on the copies of $H_{k-1}$ with intervals $[i,i+1)$. In the other case, we would consider the copies of $H_{k-1}$ with intervals $[d+1-i,d+1-i+1)$. We only need the interval weights to be increasing right-to-left. Note that for the largest index $\hat{i}=\max(I_{2})$ all vertices $a_{i}$ with $i\in I_{2}-\hat{i}$ lie below the edge $a_{\hat{i}}v^{\prime}$.
Now consider for each $i\in I_{2}-\hat{i}$ the copy of $2$-tree $H_{k-1}$ with root edge $b^{\prime}_{i}a_{i}$ and interval $[i,i+1)$. By induction, for each $i\in I_{2}-\hat{i}$, there is a $(k-1)$-inversion whose right end is $a_{i}$ and all of whose edge-weights lie in the interval $[i,i+1)$.
Now, if for at least one such $i$, the left end of the corresponding $(k-1)$-inversion $(e_{1},\ldots,e_{k-1})$ lies right of $a_{\hat{i}}$, then $(a_{\hat{i}}v^{\prime},e_{1},\ldots,e_{k-1})$ is the desired $k$-inversion, and we are done. Otherwise, each of the first edges from the $k-1$ corresponding $(k-1)$-inversions has its left endpoint to the left of $a_{\hat{i}}$. But then, taking edge $a_{\hat{i}}v^{\prime}$ and these $k-1$ first edges in order of increasing weights again gives a desired $k$-inversion.
To finish the proof of the theorem, it is now enough to take $\langle H_{p},w_{p}\rangle$ from 3 with any interval $[x,y)$, and then apply 1 to obtain the desired edge-weighted $2$-tree $\langle G_{p},w_{p}\rangle$ with $\operatorname{pqn}(G_{p},w_{p})\geq p$.
Appendix D Omitted Proofs of Section 6
It remains to prove 2, which is part of the proof of Theorem 6.1.
See 2
{claimproof}
Refer to Figure 6 for an illustration of our construction. We first show inductively that the edges on the left side are assigned to the same priority queues as their (adjacent) synchronization edges, and then we show the same for the edges on the right side.
Let $a_{1}$ be the edge with the leftmost left endpoint corresponding to a vertex $a\in S$. The $k-1$ heavy edges below $a_{1}$ require $k-1$ priority queues, none of which can be used by $a_{1}$ or its synchronization edge since both form a forbidden pattern with each of the heavy edges. Therefore, $a_{1}$ and its synchronization edge are assigned to the only remaining priority queue.
Let $b_{1}$ be the edge with the second leftmost left endpoint corresponding to a vertex $b\in S$. The $k-2$ heavy edges below $b_{1}$ require $k-2$ priority queues, none of which can be used by $b_{1}$ or its synchronization edge. Moreover, $a_{1}$ and its synchronization edge occupy another priority queue. Hence, only one priority queue is left for $b_{1}$ and its synchronization edge. This argument repeats for all subsequent vertices of $S$.
Now consider the edges on the right side belonging to vertices of $S$. Here, the symmetric argument applies from right to left. Hence, the edges $a_{2},b_{2},\dots$ are assigned to the same priority queue as their synchronization edges, which in turn implies that each pair of edges corresponding to a vertex from $S$ is assigned to the same priority queue.